Parking in a student lot cost $2 for the first half hour and $1.25 for each hour thereafter. A partial hour is considered the same as a full hour. What is the longest time that a student can park in this lot for $10?
The longest time is?

Answer by checkley79(3341)

2+1.25x<=10
1.25<=10-2
1.25<=8
x<=8/1.25
x<=6.4
x must be 6 or less.
Proof:
2+1.25*6<=10
2+7.5<=10
9.5<10